OpenStudy (anonymous):

shucks!!! 2r+3s=1: solve for R, I got 1-3s/2 (correct or not)

OpenStudy (anonymous):

no r=1/2-3s/2

OpenStudy (anonymous):

you must divide each term by 2 not just the 3s

OpenStudy (anonymous):

r=3/2s+1/2

OpenStudy (anonymous):

or you could use parentheses \[r=\frac{1-3s}{2}\]

OpenStudy (anonymous):

srry it should be -3/2s+1/2=r